# Real-time Monocular Full-body Capture in World Space via Sequential Proxy-to-Motion Learning

### v1

@@ Abstract
Learning-based approaches to monocular motion capture have recently shown promising results by learning to regress in a data-driven manner. However, due to the challenges in data collection and network designs, it remains challenging for existing solutions to achieve real-time full-body capture while being accurate in world space. In this work, we contribute a sequential proxy-to-motion learning scheme together with a proxy dataset of 2D skeleton sequences and 3D rotational motions in world space. Such proxy data enables us to build a learning-based network with accurate full-body supervision while also mitigating the generalization issues. For more accurate and physically plausible predictions, a contact-aware neural motion descent module is proposed in our network so that it can be aware of foot-ground contact and motion misalignment with the proxy observations. Additionally, we share the body-hand context information in our network for more compatible wrist poses recovery with the full-body model. With the proposed learning-based solution, we demonstrate the first real-time monocular full-body capture system with plausible foot-ground contact in world space. More video results can be found at our project page: https://liuyebin.com/proxycap.
11footnotetext: Corresponding authors.
@@ 1 Introduction
Figure 1: We demonstrate that our real-time monocular full-body capture system can produce accurate human motions with plausible foot-ground contact in world space.
@@
Capturing full-body motions from monocular videos is an essential technology for various applications such as gaming, VR/AR, sports analysis, etc. One ultimate goal is to achieve real-time full-body capture while being accurate and physically plausible in world space. Despite the recent advancements, this task is still far from being solved.
Compared with the optimization-based methods [49, 11, 3, 22, 13, 40, 60], learning-based approaches [16, 21, 69, 68] can predict 3D body motions directly, which are pretty efficient and have become increasingly popular. As data-driven solutions, the accuracy of learning-based approaches is bounded by the quality of supervision, i.e., the motion annotations. Meanwhile, their robustness heavily relies on the diversity of training data. Since the motion capture datasets [14, 30, 50, 12] are typically created with marker-based or multi-view systems and are limited in the diversity of body appearances and backgrounds, numerous efforts have been dedicated to enriching such data by generating pseudo labels [21, 15, 33] for in-the-wild datasets. Though these pseudo labels become increasingly aligned with the images, they lack accurate global motions due to the simplified camera assumptions and depth ambiguity. To tackle this limitation, researchers have also created synthetic data [39, 2] by rendering human models with controllable cameras. Despite the accurate motions and camera parameters of such synthetic data, there are distinct domain gaps between the real-world images and the rendered ones.
For physically plausible motion capture, there have been several attempts [44, 62] to correct the motion results in the world space via physics-based trajectory optimization [62]. However, these solutions typically require intensive computational resources, making them unsuitable for real-time applications. Recent state-of-the-art solutions [63, 27] alleviate this issue by incorporating the physics simulation in the learning process via reinforcement learning. Despite their effectiveness, these methods typically require the modeling of a known 3D scene [63, 27].
To handle the above challenges, we propose a learning-based full-body motion capture method with real-time performance and plausible physical contact with the ground plane. To this end, we start from the training data and propose a simple yet effective data generation strategy. Our key insight is to leverage the existing large-scale motion sequence data, i.e., AMASS [29], and synthesize sequential proxy-to-motion pairs with virtual camera settings. More specifically, we adopt the 2D skeleton sequences as proxy representations and associate them with the 3D rotational motion sequences in world space. Such a data generation strategy addresses the limitation of previous datasets in the aspects of scale, accuracy, and generalization: i) the data can be synthesized at a large scale and high quality as both body motions and virtual camera settings are accurate in world space; ii) as one of the well-studied representations, 2D skeletons can be estimated in a pretty accurate and robust manner for existing state-of-the-art toolkits, which mitigates the generalization issue of previous synthetic datasets; iii) The proxy data can be synthesized individually and then integrated to form a full-body proxy dataset by combining hand gestures and facial expressions with body motions. Moreover, the proxy sequences also help handle the depth ambiguity compared to previous work [51] using the proxy representations of a single frame.
Though the proposed proxy dataset can be more accurate and large-scale, it remains challenging for regression-based networks to learn physically plausible motions from proxy data. To address this, we propose a network architecture that takes a proxy sequence as input and produces 3D motions in world space. Our network first predicts coarse 3D motions and then refines them to be more accurate and physically plausible. Specifically, we introduce a contact-aware neural motion descent module for iterative motion refinement. In this module, we explore a cross-attention layer by using the contact status, proxy misalignment, and the current motion state as the query, key, and value. By doing so, the neural descent module can produce more accurate motion updates based on foot-ground contact and motion misalignment with the proxy observations.
Our solution is well-suited for full-body motion capture by leveraging the full-body proxy dataset. For better body-hand compatibility, our network further incorporates the context information from the hand part to predict more compatible wrist poses in the full-body model. As a learning-based solution, our method for full-body motion prediction can be trained end-to-end and run in real-time during inference. Our contributions can be summarized as follows:
  • To overcome the accuracy and generalization issues, we adopt 2D skeleton sequences as proxy representations and synthesize proxy data from accurate 3D rotational motions in world space. The proxy data of different body parts can also be integrated as one with full-body motions.
@@
  • For accurate and physically plausible predictions, we propose a contact-aware neural motion descent module so that our network can be aware of foot-ground contact and motion misalignment with the proxy observations.
  • For full-body motion capture, we leverage the body-hand context information in our network so that the wrist poses can be more compatible with the full-body model.
@@
  • We demonstrate a real-time monocular full-body capture with plausible foot-ground contact in world space.
@@ 2 Related Work
Monocular motion capture has been an active research field in recent years. We give a brief review of the works related to ours and refer readers to [54] for a more comprehensive survey.
@@
Neural Decent for Motion Capture. Traditional optimization-based approaches [3] typically fit 3D parametric models to the 2D evidence but suffer from initialization sensitivity and the failure to handle challenging poses. To achieve more efficient and robust motion prediction, there are several attempts to leverage the learning power of neural networks for iterative refinement. HUND [64] proposes a learning-to-learn approach based on recurrent networks to regress the updates of the model parameters. Song et al. [51] propose the learned gradient descent to refine the poses of the predicted body model. Similar refinement strategies are also exploited in PyMAF [69] and LVD [7] by leveraging image features as inputs. In our work, we propose a contact-aware neural decent module and exploit the usage of cross-attention for more effective motion updates.
Physical Plausibility of Motion Capture. Though existing monocular motion capture methods can produce well-aligned results, they may still suffer from the artifacts such as ground penetration and foot skating. For more physically plausible reconstruction, previous works [20, 62] have made attempts to leverage more accurate camera models during the learning process. To encourage proper contact of human meshes, Rempe et al. [45] propose a physics-based trajectory optimization to learn the body contact information explicitly. HuMoR [44] introduces a conditional VAE to learn a distribution of pose changes in the motion sequence, providing a motion prior for more plausible human pose prediction. LEMO [71] learns the proposed motion smoothness prior and optimizes with the physics-inspired contact friction term. Despite their plausible results, these methods typically require high computation costs and are unsuitable for real-time applications. For more effective learning of the physical constraints, there are several attempts [63, 27] to incorporate the physics simulation in the learning process via reinforcement learning. However, these methods [63, 27] typically depend on 3D scene modeling due to the physics-based formulation. In our work, we achieve real-time capture with plausible foot-ground contact in world space by designing novel networks and leveraging accurate motion supervision from the proxy data.
@@ 3 Proxy Data Generation
To address the precision and generalization challenges associated with extant datasets, we leverage 2D skeleton sequences as proxy representations and synthesize sequential proxy-to-motion data utilizing precise 3D rotational motions in world space. In this section, we begin by presenting the camera and motion decoupling approach in Section 3.1. Subsequently, we introduce the synthetic data generation process in Section 3.2.
3.1 Camera and Motion Decoupling
The traditional world-to-image projection typically involves transferring the world coordinates of the object to the camera space and then projecting them to the image plane. To simplify the process, previous methods [42, 74, 10, 16, 21, 69] directly regress human poses in the camera coordinate system and feed them to a projection function. However, such a simplified process introduces an ambiguous problem applying the same projection function to different scenes. Motivated by previous work [20, 66], we address this issue by decomposing the camera pose and the human motion, which allows us to predict the camera pose and the human motion in world space and calculate the foot-ground contact more directly.
In the following, we adopt the classical pinhole camera model to describe our camera settings. As shown in Fig. 3, the camera is fixed along the z-axis and positioned at the height of $h$ above the ground. The camera orientation is set to face squarely towards the front to reduce the degree of camera pose freedom with only two parameters of pitch $\alpha$ and roll $\phi$. In this setting, the camera rotation $R_{c}$ and transition $T_{c}$ can be calculated as following:
$\begin{cases}R_{c}=euler2mat(\alpha,\phi,0)\\ T_{c}=-R_{c}^{\top}\cdot(0,0,h)^{\top}\quad x\leq 0.\end{cases}$ (1)
For body motion in world space, it can be represented with the rotational parameters ${\theta}^{w}$ and global transition ${t}^{w}$. With accurate ${t}^{w}$, the body model can step on the ground $z_{w}=0$ with plausible foot-ground contact. Let the parameters of body motions in camera space be ${\theta}^{c},{t}^{c}$, their relationship with the world-space motion can be written as $\mathcal{M}(\theta^{c},t^{c})=R_{c}\cdot\mathcal{M}(\theta^{w},t^{w})+T_{c}$, where $\mathcal{M}(\theta,t)$ denotes the body model.
Figure 2: Illustration of the proposed method. Our method takes the estimated 2D skeleton poses of the body and hand as inputs and recovers the 3D motions with plausible foot-ground contact in world space.
Figure 3: Illustration of the camera pose and human motion decoupling in world space to obtain physical priors.
3.2 Data Synthesis and Integration
In our method, we adopt the 2D skeleton sequences as the proxy representation. To generate proxy-to-motion pairs with accurate human kinematics, we synthesize proxy sequences based on the existing motion datasets with virtual cameras. In the following, we describe the synthesis and integration of different types of labels in our proxy data, including body motions, hand gestures, and contact labels.
Body proxy data. We adopt the motion sequences from the AMASS dataset [29] to generate proxy-to-motion pairs for the body part. The AMASS dataset is a large-scale body motion sequence dataset that comprises 3772 minutes of motion sequence data, featuring diverse and complex body poses. We downsample the motion data to 60 frames per second, resulting in 9407K frames.
@@ 3.2 Data Synthesis and Integration
Integration with contact label. To generate the contact label, we follow LEMO [70] and select 6 vertices on the foot of the SMPL and calculate the continuous contact indicators as follows:
@@
where $v_{i}$ and $z_{i}$ denote the velocity in XY-plane and the height to the ground of the selected vertex. $v_{max}$ and $z_{max}$ is setting to $0.2m/s$ and $0.08m$, the hyper-parameter $k_{v}$ and $k_{z}$ is set to $0.04$ and $0.008$. The resulting contact label can be directly integrated into the proxy data.
Camera setting. We adopt the same camera setup used in existing public datasets [14, 58] for virtual camera settings. Specifically, we set the fov from $50^{\circ}$ to $70^{\circ}$ and sampled the camera pose from: $\alpha\sim\mathcal{U}(0^{\circ},30^{\circ})$, $\phi\sim\mathcal{U}(-5^{\circ},5^{\circ})$, to ensure variability in camera viewpoints. Meanwhile, we will apply a random rotation $\gamma_{aug}\sim\mathcal{U}(0^{\circ},360^{\circ})$ around the z-axis to human and apply a random y-axis displacement $d_{aug}\sim\mathcal{U}(2.5m,6m)$ to simulate all cases that the person is near or far to the camera to augment the data. We generate pseudo 2D skeleton annotations by adding Gaussian noise $\Delta X\sim\mathcal{N}(0,0.003m)$ to the 3D joints of the body model and then project them to the image plane.
@@ 4 Method
In our approach, the skeleton poses are first detected from a monocular video and then fed into our network to produce 3D motions in world space, as illustrated in Fig. 2. In the following, we describe the proposed sequential proxy-to-motion learning scheme to achieve accurate and physically plausible motion prediction in world space.
4.1 Sequential Full-body Motion Recovery
At the first stage of the proposed scheme, the skeleton sequences are processed by temporal encoders and then fed into regressors for the predictions of full-body motion in world space. Specifically, the temporal encoders are used to extract features from sequential 2D skeletons of the body and hand parts. Following previous baseline [42], we build these encoders upon temporal dilated convolutional networks. Moreover, inspired by previous skeleton-based 3D human pose estimation [31, 42, 74, 10, 66], we adopt two separate branches in the body-specific regressors to learn the global trajectory and root-relative body motions. Specifically, the trajectory branch learns the camera parameters $\alpha,\phi,h$ and the global human transition $t_{b}$, while the motion branch learns the shape parameter $\beta_{b}$ and pose parameter $\theta_{b}$ of the body model. Separating these two prediction tasks helps to decouple the global and local motions and hence produce more accurate motions in world space.

### v3

# ProxyCap: Real-time Monocular Full-body Capture in World Space via Human-Centric Proxy-to-Motion Learning

@@ Abstract
Learning-based approaches to monocular motion capture have recently shown promising results by learning to regress in a data-driven manner. However, due to the challenges in data collection and network designs, it remains challenging for existing solutions to achieve real-time full-body capture while being accurate in world space. In this work, we introduce ProxyCap, a human-centric proxy-to-motion learning scheme to learn world-space motions from a proxy dataset of 2D skeleton sequences and 3D rotational motions. Such proxy data enables us to build a learning-based network with accurate world-space supervision while also mitigating the generalization issues. For more accurate and physically plausible predictions in world space, our network is designed to learn human motions from a human-centric perspective, which enables the understanding of the same motion captured with different camera trajectories. Moreover, a contact-aware neural motion descent module is proposed in our network so that it can be aware of foot-ground contact and motion misalignment with the proxy observations. With the proposed learning-based solution, we demonstrate the first real-time monocular full-body capture system with plausible foot-ground contact in world space even using hand-held moving cameras. Our project page is https://zhangyux15.github.io/ProxyCapV2.
Figure 1: The proposed method, ProxyCap, is a real-time monocular full-body capture solution to produce accurate human motions with plausible foot-ground contact in world space.
@@ 1 Introduction
Motion capture from monocular videos is an essential technology for various applications such as gaming, VR/AR, sports analysis, etc. One ultimate goal is to achieve real-time capture while being accurate and physically plausible in world space. Despite the recent advancements, this task is still far from being solved, especially under the settings of in-the-wild captures with hand-held moving cameras.
Compared to optimization-based methods [50, 10, 4, 20, 12, 40, 64], learning-based approaches [15, 19, 70, 72] can directly regress human poses from images, significantly enhancing computational efficiency while addressing the inherent issues in optimization-based methods of initialization sensitivity and local optima entrapment. As data-driven solutions, the performance and generalization capabilities of learning-based methodologies are heavily bounded by the accuracy and diversity of the training data. Unfortunately, existing datasets are unable to simultaneously meet these requirements. On the one hand, datasets with sequential ground truth 3D pose annotations  [13, 30, 51, 11] are mostly captured by marker-based or multi-view systems, which makes it hard to scale up to a satisfactory level of diversity in human appearances and backgrounds. On the other hand, numerous in-the-wild datasets [1, 23] excel in the richness of human and scenario diversity but they lack real-world 3D motions and most of them only provide individual images instead of videos. Recently, researchers tried to create synthetic data [38, 2, 3] by rendering human avatars with controllable cameras, but it remains difficult to bridge domain gaps between the real-world images and the rendered ones, and is too expensive to scale up.
In this paper, we follow the spirit of creating synthetic data, but turn to render 2D proxy representations instead of person images. By using proxy representations (i.e., silhouettes [39, 60], segmentations [16, 37, 47, 16], IUV [71, 62] and 2D skeletons [39, 54, 58, 31, 32, 52, 28, 21]), the whole motion capture pipeline can be divided into two steps: image-to-proxy extraction and proxy-to-motion lifting. In the divided pipeline, the image-to-proxy extraction is pretty accurate and robust as there are plenty of annotated 2D ground truths in real-world datasets, while the proxy-to-motion step can leverage more diverse training data to mitigate the generalization issue and reduce the domain gap. Here we adopt the 2D skeleton sequences as the proxy representation for its simplicity and high correlation with 3D motion. Meanwhile, we combine random virtual camera trajectories upon the existing large-scale motion sequence database, i.e., AMASS [29], to synthesize nearly infinite data for learning proxy-to-motion lifting.
Though the proposed proxy dataset can be generated at large scales, there remain two challenges for regression-based networks to learn physically plausible motions from proxy data: how to recover i) world-space human motion under moving cameras, and ii) physically plausible motion with steady body-ground contact. For a video captured with moving cameras, the trajectories/rotations of humans and cameras are coupled with each other, making the recovery of world-space human motion extremely difficult. To address this issue, the latest solutions [18, 63] estimate the camera poses from the background using SfM [56, 48], and then estimate the human motion from a camera-centric perspective. However, the SfM requires texture-rich backgrounds and may fail when the foreground moving character dominates the image. Their post-processing optimization pipelines are also not suitable for real-time applications due to expensive computational costs. Besides, these previous solutions learn human motion in a camera-centric perspective like [17], which is actually ambiguous for the regression network.
In this paper, we would point out that one of the main challenges arises from the camera-centric settings in previous solutions. In such a setting, the same motion sequence captured under different cameras will be represented as multiple human motion trajectories, making it difficult for the network to understand the inherent motion prior. In contrast, we propose to learn human-centric motions to ensure consistent human motion outputs under different camera trajectories in synthetic data. Specifically, our network learns the local translations and poses in a human coordinate system, together with the relative camera extrinsic parameters in this space. After that, we accumulate the local translations of humans in each frame to obtain the global camera trajectories. Benefiting from the proposed proxy-to-motion dataset, we are able to synthesize different camera trajectories upon the same motion sequence to learn the human motion consistency. In this way, our network can disentangle the human poses from the moving camera more easily via the strong motion prior without SfM.
On top of human-centric motion regression, we further enhance the physical plausibility by introducing a contact-aware neural motion descent module. Specifically, our network first predicts coarse motions and then refines them iteratively based on foot-ground contact and motion misalignment with the proxy observations. Compared with the global post-processing optimization used in previous work [67, 18, 63], our method learns the descent direction and step instead of explicit gradient back-propagation. We demonstrate that our method, termed ProxyCap, is more robust and significantly faster to support real-time applications. To sum up, our contributions can be listed as follows:
  • To tackle the data scarcity issue, we adopt 2D skeleton sequences as proxy representations and generate proxy data in world space with random virtual camera trajectories.
  • We design a network to learn motions from a human-centric perceptive, which enables our regressor to understand the consistency of human motions under different camera trajectories.
@@
  • We further propose a contact-aware neural descent module for more accurate and physically plausible predictions. Our network can be aware of foot-ground contact and motion misalignment with the proxy observations.
@@
  • We demonstrate a real-time full-body capture system with plausible body-ground contact in world space under moving cameras.
@@ 2 Related Work
Monocular motion capture has been an active research field recently. We give a brief review of the works related to ours and refer readers to [57] for a more comprehensive survey.
@@
Neural Decent for Motion Capture. Traditional optimization-based approaches [4] typically fit 3D parametric models to the 2D evidence but suffer from initialization sensitivity and the failure to handle challenging poses. To achieve more efficient and robust motion prediction, there are several attempts to leverage the learning power of neural networks for iterative refinement. HUND [68] proposes a learning-to-learn approach based on recurrent networks to regress the updates of the model parameters. Song et al. [52] propose the learned gradient descent to refine the poses of the predicted body model. Similar refinement strategies are also exploited in PyMAF [70] and LVD [6] by leveraging image features as inputs. In our work, we propose a contact-aware neural decent module and exploit the usage for more effective motion updates.
Plausible Motion Capture in World Space. Though existing monocular motion capture methods can produce well-aligned results, they may still suffer from artifacts such as ground penetration and foot skating in world space. For more physically plausible reconstruction, previous works [17, 67] have made attempts to leverage more accurate camera models during the learning process. To encourage proper contact of human meshes, Rempe et al. [44] propose a physics-based trajectory optimization to learn the body contact information explicitly. HuMoR [45] introduces a conditional VAE to learn a distribution of pose changes in the motion sequence, providing a motion prior for more plausible human pose prediction. LEMO [73] learns the proposed motion smoothness prior and optimizes with the physics-inspired contact friction term. Despite their plausible results, these methods typically require high computation costs and are unsuitable for real-time applications. For more effective learning of the physical constraints, there are several attempts [66, 27] to incorporate the physics simulation in the learning process via reinforcement learning. However, these methods [66, 27] typically depend on 3D scene modeling due to the physics-based formulation. Recently, there are also attempts to recover camera motion via SLAM techniques [63, 18] or regress the human motion trajectory [43, 55]. Despite the progress, it remains challenging for these methods to run in real-time or produce physically plausible in world space. In our work, we achieve real-time capture with plausible foot-ground contact in world space by designing novel networks to learn human-centric motion.
@@ 3 Proxy Data Generation
Figure 2: Illustration of the proposed method ProxyCap. Our method takes the estimated 2D skeletons from a sliding window as inputs and estimates the relative 3D motions in the human coordinate space. These local movements are accumulated frame by frame to recover the global 3D motions. For more accurate and physically plausible results, a contact-aware neural motion descent module is proposed to refine the initial motion predictions.
To tackle the data issue, we synthesize sequential proxy-to-motion data based on 2D skeletons and their corresponding 3D rotational motions in world space. In the following, we describe the synthesis and integration of different types of labels in our proxy data, including body motions, hand gestures, and contact labels.
Body proxy data. We adopt the motion sequences from the AMASS dataset [29] to generate proxy-to-motion pairs for the body part. The AMASS dataset is a large-scale body motion sequence dataset that comprises 3,772 minutes of motion sequence data, featuring diverse and complex body poses. We downsample the motion data to 60 frames per second, resulting in 9407K frames.
@@
Integration with contact labels. We calculate the continuous contact indicators $ind$ for 3D skeletons as follows:
@@
where $v_{i}$ and $z_{i}$ denote the velocity and the height to the xz-plane of the given joint. $v_{max}$ and $z_{max}$ is set to $0.2m/s$ and $0.08m$, and $k_{v}$ and $k_{z}$ is set to $0.04$ and $0.008$.
Camera setting. For each 3D motion sequence, we generate 2D proxy data under different virtual camera trajectories (four cameras in practice). Such proxy data enhances the learning of the inherent relationship between 2D proxy and 3D motions and the consistency across different camera viewpoints.
Specifically, we uniformly sample the field of view (FOV) from $30^{\circ}$ to $90^{\circ}$ for the intrinsic parameters of cameras. When setting the extrinsic parameters for camera trajectories, we position the cameras at distances ranging from $1$ meter to $5$ meters around the human, and at heights varying from $0.5$ meters to $2$ meters to the ground. Finally, we generate pseudo 2D skeleton annotations by projecting 3D joints into the 2D pixel-plane using these virtual cameras. Moreover, to simulate the jitter of 2D detectors, we add Gaussian noise $\Delta X\sim\mathcal{N}(0,0.01)$ to 3D joints before projections.
@@ 4 Method
As illustrated in Fig. 2, we first detect the 2D skeletons from images and input them into our proxy-to-motion network to recover 3D local motions in human space. These relative local motions are transformed into a world coordinate system and accumulated along the sliding window. Additionally, we leverage a neural descent module to refine the accuracy and physical plausibility. In this section, we start with introducing the human-centric setting of our motion prediction.
4.1 Human-Centric Motion Modeling
For more accurate camera modeling, we adopt the classical pinhole camera model instead of using the simplified orthogonal or weak perspective projection [41, 77, 9, 15, 19, 70]. As shown in Fig. 3, we transform the global human motion and the camera trajectories into local human space from two adjacent frames, where we adopt $\{\beta\in\mathbf{R^{10}},\theta\in\mathbf{R^{22\times 3}},t\in\mathbf{R^{3}}\}_{t}$ to denote the parameters of shape, pose, and translation at frame $t$, and $\{R\in\mathbf{R^{3\times 3}},T\in\mathbf{R^{3}}\}_{t}$ to denote the camera extrinsic parameters. Given the pose and shape parameters, the joints and vertices can be obtained within the SMPL Layer: $\{J,V\}=\mathcal{X}\left(\beta,t,\theta,g\right)$. In the following, we use the subscript $H$ and $W$ to distinguish between the human coordinate system and the world coordinate system.
During the learning of human-centric motions, we adopt a similar setting used in previous works on temporal human motion priors [24, 45, 67]. Specifically, each motion sequence will be normalized by a rigid transformation to remove the initial translation in x-z plane and rotate itself with the root orientation heading to the z-axis direction. With this setting, our network can learn the relative movements of humans, which are independent of observation viewpoints. The detailed implementation of the human-to-world coordinate transformation and the global motion accumulation in world space can be found in our supplementary material.
Figure 3: Illustration of decoupling the world-space motion into the human-centric coordinates and relative camera poses.
4.2 Sequential Full-body Motion Initialization
As shown in Fig. 2, at the first stage of our method, the skeleton sequences are processed by temporal encoders and then fed into decoders for the predictions of the initial motion $\{\mathbf{\beta}^{0},\mathbf{\theta}^{0},\mathbf{t}^{0}\}$, initial camera $\{\mathbf{R}^{0}_{\mathbf{H}},\mathbf{T}^{0}_{\mathbf{H}}\}$, the contact indicator $\mathbf{ind}$, and the hand poses $\mathbf{g}$. Following previous baseline [41], we build these encoders upon temporal dilated convolution networks.